Consider a random variable Y = 2X + 7, where X is a random variable with mean 3 and standard deviation 2.
What is Var(Y)?

Answer by Edwin McCravy(20059)

If X is a random variable with mean μ and standard deviation σ,
then the random variable Y = aX + b has mean = aμ + b, and 
standard deviation aσ.

Therefore, since X is a random variable with mean 3 and standard deviation 2,
random variable Y = 2X + 7 has mean 2(3) + 7 = 13 and standard deviation 
2(2) = 4, so its variance is 42 = 16.

Edwin